OpenStudy (anonymous):

graph f(x)=2/x-3 -2 domain of graph range of graph

OpenStudy (anonymous):

this function? \[f(x) = \frac{2}{x-3} - 2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, lets graph it first. we should think about x and y axis intecepts, first for y intercept, find the value of y when x = 0 can you do that for me?

OpenStudy (anonymous):

y = f(x)

OpenStudy (anonymous):

so what is f(0) ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

\[f(x) = \frac{2}{x-3} -2\] \[f(0) = \frac{2}{0-3} -2 = \frac{2}{-3} - 2 = -\frac{8}{3}\] ok?

OpenStudy (anonymous):

okay i understand that part

OpenStudy (anonymous):

so this tells us that when x = 0 , y = -8/3 :  next we have to find the x value when y = 0

OpenStudy (anonymous):

\[0 = \frac{2}{x-3} - 2\] \[2 = \frac{2}{x-3}\] \[2(x-3) = 2\] \[(x-3) = 1\] \[x = 4\]